# Euclid Quick Data Release (Q1)

### v1

@@ 4.4 Comparison between Euclid and external cluster observables
The SZ signal is detected in the three richness bins. We also present the results for two subsamples: the matched subsample (blue triangles) containing the 110 detections matched with catalogues and meta-catalogues corresponding to numbers in brackets in Table 1; and the unmatched subsample (black squares) containing the remaining detections. The SZ signal is detected in the highest richness (third) bin for the two subsamples with the same amplitude, but is detected/not detected in the two lowest richness (first and second) bins for the matched/unmatched subsample respectively. In particular, the difference in SZ flux between the matched and unmatched detections in the first/second bin is respectively, indicating that the unmatched subsample contains significantly less hot gas than the matched subsample for richness . Investigations with deeper SZ observations (e.g., ACT or SPT-3G) and deep X-ray observations (e.g., XMM or eROSITA) will be necessary to quantify better the gas content of the newly discovered Euclid detections.
This difference in the gas content of the matched and unmatched subsample with may be due to the fact that a large fraction of our external catalogues and meta-catalogues are SZ or X-ray selected, so most probably contain gas rich systems. The unmatched subsample is thus expected to contain more poor gas systems. It is also possible that projection effects in the Euclid detection are boosting richness for intrinsically poor systems, whose mass is consistent with a very low SZ and X-ray signal. While this effect is expected to be marginal for the richest systems (for which SZ flux is comparable for matched and unmatched), it could be significant at lower richness. This result should also be compared with the result from Euclid Collaboration: Gouin et al. (2025), in which they measured the connectivity of the eROSITA clusters for gas rich and gas poor systems, finding that gas poor systems are more connected to the cosmic web than gas rich systems. Similar findings are presented in Popesso et al. (2024), whereby optically-detected (and X-ray-undetected) systems are predominantly situated in filaments, while the X-ray detected ones are predominantly located in nodes.
@@ 4.5 Unmatched detections from external data sets
Fig. 14 shows the distribution in SZ-based total mass, RedMaPPer richness, and eROSITA X-ray based mass estimate, for the matched and unmatched clusters falling within the EDF-S, where the expected coverage is above 80% (see Sect. 2.2). We see a consistent trend across optical, X-ray and millimetre wavelengths, whereby the unmatched populations systematically occupy the lower richness and lower mass ends of the overall distribution.
Since the joint Euclid catalogue is cut at high S/N by construction, it does not completely sample the underlying cluster population, particularly at lower richnesses (78% of the unmatched RedMaPPer objects have a richness less than 30). This reinforces the hypothesis that incompleteness in the cluster sample at this stage is largely driven by the high S/N threshold on the individual detection catalogues, in addition to known issues of incompleteness in the Euclid galaxy catalogues at lower redshifts, described in more detail in Klein et al. (in preparation).
@@ 4.6 The potential for new Euclid cluster detections
For the consideration of new, potentially undetected clusters in the Q1 fields, we expanded our search beyond the list of catalogues and meta-catalogues described in Sect. 4.1 for the redshift validation and measurement of positional scatter. We performed a cross-match of all the detections within the NASA/IPAC Extragalactic Database (NED) 999NED is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology., using a matching radius of two arcminutes, and a redshift constraint of . We also perform a match to the Massive and Distant Clusters of the Wise Survey 2 (MaDCoWS2), described in Thongkham et al. (2024) and Wen and Han 2024 (WH24) cluster catalogue (Wen & Han 2024), using an identical scheme to the one described above, simply replacing the NED redshifts with those from MaDCoWS2/WH24.
@@ 4.7 Beyond : the case for Euclid
While we impose an upper redshift limit on our joint cluster sample for this first analysis due to the limitations in photometric redshift estimation described in Sect. 2.3, the detection algorithms as well as the richness estimation are nevertheless applied on the nominal Euclid redshift range, which extends to a redshift of . To this end, we study the number of joint clusters which fall in the redshift range to evaluate their reliability as a means to highlight one of the key scientific goals of Euclid– the detection and characterisation of the high redshift cluster population, as Sartoris et al. (2016) showed that formally half of the cosmological constraining power of the Euclid cluster survey comes from clusters in the redshift range . In total, we find 58 systems across the three fields (four in the EDF-F, 10 in the EDF-S, and 44 in the EDF-N). The asymmetry in the number of detections between the north and south is driven by the apparent higher rate of contamination in the north in this field, in part due to its higher stellar density. We use the same criteria as described in Sect. 4.6 to determine the reliability of these high redshift systems, noting 15 in total, which display clear visual overdensities in Euclid and external imaging data where applicable. The complete properties, including positions, redshifts and richness estimates for these 15 systems are shown in Table 3. We display postage stamps of six examples of these candidates across the three fields in Fig. 15.
@@ 5 External spectroscopic validation of clusters
The spectroscopic redshift of the cluster was estimated using the biweight technique which has been shown to be appropriate for a small number of objects (Beers et al. 1990). After identifying the peak in redshift space, the criteria in photometric redshift was relaxed, and the radius was increased to 2 Mpc in order to improve statistics. Galaxies with a radial velocity offset larger than in the cluster rest frame were rejected. The estimates of the spectroscopic redshift, its error, and the final number of galaxies used within 2 Mpc are provided in Table 2 for those clusters with more than 3 redshifts remaining at the end of the procedure.
@@ 6 Conclusions
We also foresee improvements in contamination rejection and masking prior to DR1, although both of these concerns are already manageable with the Q1 data set. Because we expect that the external band photometry will improve significantly following the observation of more fields to better calibrate OU-PHZ templates, in addition to bias corrections in the photometric pipeline, we can expect that the results will improve considerably for DR1, whilst already demonstrating a successful application of multiple aspects of the Euclid cluster workflow for the detection and characterisation of optically-selected galaxy clusters.
@@ Acknowledgements.
- Bulbul et al. (2024) Bulbul, E., Liu, A., Kluge, M., et al. 2024, arXiv e-prints, arXiv:2402.08452
@@
- Costanzi et al. (2020) Costanzi, M., Saro, A., Bocquet, S., et al. 2020, arXiv preprint arXiv:2010.13800
@@
- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted, arXiv:2503.15302
- Euclid Collaboration: Bergamini et al. (2025) Euclid Collaboration: Bergamini, P., Meneghetti, M., Acebron, A., et al. 2025, A&A, submitted, arXiv:2503.15330
- Euclid Collaboration: Castander et al. (2024) Euclid Collaboration: Castander, F., Fosalba, P., Stadel, J., et al. 2024, A&A, accepted, arXiv:2405.13495
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
- Euclid Collaboration: Enia et al. (2025) Euclid Collaboration: Enia, A., Pozzetti, L., Bolzonella, M., et al. 2025, A&A, submitted, arXiv:2503.15314
@@
- Euclid Collaboration: Gouin et al. (2025) Euclid Collaboration: Gouin, C., Laigle, C., Sarron, F., et al. 2025, A&A, submitted, arXiv:2503.15332
- Euclid Collaboration: Ingoglia et al. (2024) Euclid Collaboration: Ingoglia, L., Sereno, M., Farrens, S., et al. 2024, A&A, accepted, arXiv:2409.02783
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, accepted, arXiv:2405.13493
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H. J., Benson, K., Dolding, C., et al. 2025, A&A, submitted, arXiv:2503.15303
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
- Euclid Collaboration: Ragagnin et al. (2024) Euclid Collaboration: Ragagnin, A., Saro, A., Andreon, S., et al. 2024, A&A, accepted, arXiv:2412.00191
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted, arXiv:2503.15305
- Euclid Collaboration: Tucci et al. (2025) Euclid Collaboration: Tucci, M., Paltani, S., Hartley, W. G., et al. 2025, A&A, submitted, arXiv:2503.15306
@@
- Lesci et al. (2020) Lesci, G. F., Marulli, F., Moscardini, L., et al. 2020, arXiv e-prints, arXiv:2012.12273
@@
- Romanello et al. (2023) Romanello, M., Marulli, F., Moscardini, L., et al. 2023, arXiv preprint arXiv:2310.12224
@@
- Sadibekova et al. (2024) Sadibekova, T., Arnaud, M., Pratt, G. W., Tarrío, P., & Melin, J. B. 2024, arXiv e-prints, arXiv:2402.01538
@@
- Tarrío et al. (2019) Tarrío, P., Melin, J. B., & Arnaud, M. 2019, A&A, 626, A7
@@ Appendix B Detailed breakdown of Q1 cluster matches to external data sets
Cluster catalogue or database | Matches with Euclid Q1 | Unique matches with Euclid Q1 |
@@
40 additional matches from MaDCoWS2: MOO2 J04046-47244, MOO2 J04173-47517, MOO2 J03530-48222, MOO2 J04160-49125, MOO2 J04105-46266, MOO2 J03483-49427, MOO2 J04043-48283, MOO2 J04248-47013, MOO2 J03493-48574, MOO2 J04077-47479, MOO2 J04232-46499, MOO2 J03478-48109, MOO2 J04136-47043, MOO2 J03476-49006, MOO2 J04143-47118, MOO2 J04069-49247, MOO2 J03542-47435, MOO2 J04107-48239, MOO2 J03585-50417, MOO2 J04120-49190, MOO2 J04204-46153, MOO2 J03549-48532, MOO2 J03483-49409, MOO2 J04192-46256, MOO2 J04119-47495, MOO2 J04205-48161, MOO2 J04023-48042, MOO2 J04066-47457, MOO2 J04123-47534, MOO2 J04186-49179, MOO2 J03532-50396, MOO2 J03522-49334, MOO2 J04008-49103, MOO2 J04154-48413, MOO2 J04151-48009, MOO2 J03509-49317, MOO2 J03541-48153, MOO2 J04152-49079, MOO2 J03580-47265, MOO2 J04006-47046
51 additional matches from WH24: J040151.3-502854, J035147.6-480407, J035004.3-504126, J041229.0-484015, J034824.7-503740, J041902.0-470112, J041914.1-482655, J042540.5-480811, J041138.8-465340, J040846.4-473539, J040428.8-484850, J040200.5-463646, J040611.8-500553, J035221.9-473201, J040605.3-494610, J041855.8-475812, J040712.4-461328, J040216.1-482417, J041702.2-470125, WH-J040140.2-491201, J040721.5-464740, J035907.0-463805, J042219.3-475103, J040548.4-465317, WH-J035705.1-494202, J040109.4-463138, J034734.6-481427, J040312.3-462817, J040820.6-495307, J034221.0-492531, J041226.6-490336, J040725.5-473706, J042231.4-463027, WH-J035607.4-473906, J042432.7-465141, J040358.4-461135, J042427.8-480720, J040453.8-504000, J041013.3-465056, J035540.3-472706, J041116.5-484601, J041859.3-472158, J040128.6-504238, J034850.9-485315, J041916.6-465144, WH-J040107.5-463144, J040641.8-480447, J035647.7-482301, WH-J041325.4-493214, J040732.6-471725, J035924.7-473453
Cluster catalogue or database | Matches with Euclid Q1 | Unique matches with Euclid Q1 |
@@
23 additional matches from MaDCoWS2: MOO2 J03388-28477, MOO2 J03303-29472, MOO2 J03295-26569, MOO2 J03380-28309, MOO2 J03319-26335, MOO2 J03279-26298, MOO2 J03354-27323, MOO2 J03247-28214, MOO2 J03297-26347, MOO2 J03304-29520, MOO2 J03299-26595, MOO2 J03280-26460, MOO2 J03280-26470, MOO2 J03384-29032, MOO2 J03325-26425, MOO2 J03323-29267, MOO2 J03266-28443, MOO2 J03341-26222, MOO2 J03346-29440, MOO2 J03267-27573, MOO2 J03251-29069, MOO2 J03330-29217, MOO2 J03363-27493
29 additional matches from WH24: J032320.8-275138, J032828.4-295942, J032318.6-273811, J032552.8-263516, J032335.7-280335, J032549.5-285005, J032607.4-263627, J034011.8-290926, WH-J033919.5-291405, J033758.0-292753, J032938.6-300310, J033133.2-261821, J033436.3-273501, J033558.8-274241, J033917.3-284720, J034016.8-291829, J033602.4-293258, WH-J032904.2-285733, J033006.9-300027, J032733.5-274222, J033926.3-290414, J033543.5-292251, J033636.4-263933, J034021.0-280456, WH-J033442.8-293101, WH-J033619.7-272146, J032730.6-264837, J032521.7-271246, WH-J032529.0-264822
Cluster catalogue or database | Matches with Euclid Q1 | Unique matches with Euclid Q1 |
@@
No additional match from MaDCoWS2.
47 additional matches from WH24: WH-J174727.1+663502, J175335.3+663546, J175510.5+663351, J174245.2+675649, WH-J174802.7+660700, J180548.0+652747, J173831.1+672459, WH-J175035.9+652105, J172910.9+660121, J174719.5+652517, J182238.7+653505, J174731.5+655955, J180722.3+674743, J175403.7+641941, J175946.8+652948, J174744.3+652823, J175253.9+641430, J181642.9+654440, J175548.4+670654, J182229.6+664025, J175949.6+641218, J180503.2+660125, J175221.2+670302, WH-J174739.8+642357, J174739.1+632030, WH-J175908.9+675602, J180358.0+662736, J175342.6+633801, J174524.6+674041, J174309.1+672117, J181607.9+681651, J175100.6+673434, J175226.8+642809, J175553.0+650743, J180036.2+661951, J175020.1+633659, J175916.7+642714, J173814.6+644140, J182006.2+663324, J182244.8+670116, J173723.1+654527, J175127.6+671719, J181908.2+644601, J180309.4+655050, WH-J180457.9+641550, WH-J180356.5+654830
@@ Appendix C Spectroscopic validation with OU-SPE
The detection algorithms estimate cluster redshifts based on the mean photometric redshift of member galaxies. To confirm and further refine these estimates, we examined the Euclid spectroscopy of galaxies in the cluster fields, as provided in the Q1 data release. We limited our inspections of the Euclid spectroscopy to those candidate clusters with , the lower limit at which the H lines can be detected given the spectral sensitivity range of the Euclid red grism (Euclid Collaboration: Mellier et al. 2024). This leaves us with 100 cluster candidates (17 in the EDF-F, 43 in the EDF-S, and 40 in the EDF-N).
The Euclid OU-SPE package provides five estimates for each galaxy; in the following, we only use the highest ranked estimate – spec.spec_rank=0 in the Science Archive System (SAS)– when selecting objects for verification of their spectroscopic redshifts. We restrict our analysis to galaxies with OU-SPE redshift estimates with signal-to-noise ratio (spe_line_snr_gf in the SAS) , probability (spe_z_prob in the SAS), , with line flux to avoid spurious features, and based on more than 300 pixels (spe_npix>300). Of the galaxies passing these spectroscopic quality criteria, we only considered those within 4 arcmin from the candidate cluster center (corresponding to in the range ), and with spectroscopic redshifts within a range of , corresponding to about three times the estimated pre-launch cluster uncertainty (Euclid Collaboration: Adam et al. 2019).
In total, we selected 527 galaxies (87 in EDF-F, 197 in EDF-S, and 243 in EDF-N), and after visual inspection of their spectra we retained 170 redshift estimates as reliable (35 in EDF-F, 65 in EDF-S, and 70 in EDF-N), that is 1/3 of the total. This means we deal with, on average, less than two reliable redshifts per cluster field.

### v2

@@ 4.4 Comparison between Euclid and external cluster observables
The SZ signal is detected in the three richness bins. We also present the results for two subsamples: the matched subsample (blue triangles) containing the 110 detections matched with catalogues and meta-catalogues corresponding to numbers in brackets in footnote 9; and the unmatched subsample (black squares) containing the remaining detections. The SZ signal is detected with the same amplitude in the highest richness bin (third) for both subsamples. In the lower richness bins (first and second), the signal is detected for the matched subsample but not detected for the unmatched subsample. In particular, the difference in SZ flux between the matched and unmatched detections in the first/second bin is respectively, indicating that the unmatched subsample contains significantly less hot gas than the matched subsample for richness . Investigations with deeper SZ observations (e.g. ACT or SPT-3G) and deep X-ray observations (e.g. XMM or eROSITA) will be necessary to quantify better the gas content of the newly discovered Euclid detections.
This difference in the gas content of the matched and unmatched subsample with may be due to the fact that a large fraction of our external catalogues and meta-catalogues are SZ or X-ray selected, so most probably contain gas-rich systems. The unmatched subsample is thus expected to contain more gas-poor systems. It is also possible that projection effects in the Euclid detection are boosting richness for intrinsically poor systems, whose mass is consistent with a very low SZ and X-ray signal. While this effect is expected to be marginal for the richest systems (for which SZ flux is comparable for matched and unmatched), it could be significant at lower richness. This result should be compared with the results reported in Popesso et al. (2024), where optically detected (but X-ray undetected) systems identified with eROSITA are predominantly found in filaments, while those detected in X-rays are mainly located in nodes. Recent simulation-based studies supporting this trend are presented in Cui et al. (2024) and Marini et al. (2025).
@@ 4.5 Unmatched detections from external data sets
Fig. 15 shows the distribution in SZ-based total mass, RedMaPPer richness, and eROSITA X-ray based mass estimate, for the matched and unmatched clusters falling within the EDF-S, where the expected coverage is above 80% (see Sect. 2.2). We see a consistent trend across optical, X-ray and millimetre wavelengths, whereby the unmatched populations systematically occupy the lower richness and lower mass ends of the overall distribution, although within a limited area of approximately 27 square degrees, the statistics are sparse.
Since the joint Euclid catalogue is cut at high S/N by construction, it does not completely sample the underlying cluster population, particularly at lower richnesses (78% of the unmatched RedMaPPer objects have a richness less than 30). This reinforces the hypothesis that incompleteness in the cluster sample at this stage is largely driven by the high S/N threshold on the individual detection catalogues and the degree of intrinsic scatter shown in Fig. 4, in addition to known issues of incompleteness in the Euclid galaxy catalogues at lower redshifts, described in more detail in Klein et al. (in preparation).
@@ 4.6 The potential for new Euclid cluster detections
For the consideration of new, potentially undetected clusters in the Q1 fields, we expanded our search beyond the list of catalogues and meta-catalogues described in Sect. 4.1 for the redshift validation and measurement of positional scatter. We performed a cross-match of all the detections within the NASA/IPAC Extragalactic Database (NED) 101010NED is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology., using a matching radius of two arcminutes, and a redshift constraint of . We also perform a match to the Massive and Distant Clusters of the Wise Survey 2 (MaDCoWS2), described in Thongkham et al. (2024) and Wen and Han 2024 (WH24) cluster catalogue (Wen & Han 2024), using an identical scheme to the one described above, simply replacing the NED redshifts with those from MaDCoWS2/WH24.
@@ 4.7 Beyond : the case for Euclid
While we impose an upper redshift limit on our joint cluster sample for this first analysis due to the limitations in photometric redshift estimation described in Sect. 2.3, the detection algorithms as well as the richness estimation are nevertheless applied on the nominal Euclid redshift range, which extends to a redshift of . To this end, we study the number of joint clusters which fall in the redshift range to evaluate their reliability as a means to highlight one of the key scientific goals of Euclid– the detection and characterisation of the high redshift cluster population, as Sartoris et al. (2016) showed that formally half of the cosmological constraining power of the Euclid cluster survey comes from clusters in the redshift range . Due to the increased scatter of galaxy photometric redshifts beyond and contamination in the galaxy catalogues, we expect strong incompleteness and impurity in this range. However, we present a subsample of promising candidates that were selected based on a combination of S/N above the threshold and visual inspection. Because the purity is low at this epoch for the current catalogue, these represent a small subset of high S/N detections. We anticipate significantly improved purity in this redshift range for DR1.
In total, we find 58 systems across the three fields (four in the EDF-F, 10 in the EDF-S, and 44 in the EDF-N). The asymmetry in the number of detections between the north and south is driven by the apparent higher rate of contamination in the north in this field, in part due to its higher stellar density. We use the same criteria as described in Sect. 4.6 to determine the reliability of these high redshift systems, noting 15 in total, which display clear visual overdensities in Euclid and external imaging data where applicable. The complete properties, including positions, redshifts and richness estimates for these 15 systems are shown in Table 3. We display postage stamps of six examples of these candidates across the three fields in Fig. 16.
@@ 5 External spectroscopic validation of clusters
The spectroscopic redshift of the cluster was estimated using the biweight technique which has been shown to be appropriate for a small number of objects (Beers et al. 1990). After identifying the peak in redshift space, the criteria in photometric redshift was relaxed, and the radius was increased to 2 Mpc in order to improve statistics. Galaxies with a radial velocity offset larger than in the cluster rest frame were rejected. The estimates of the spectroscopic redshift, its error, and the final number of galaxies used within 2 Mpc are provided in Table 2 for those clusters with more than three redshifts remaining at the end of the procedure.
@@ 6 Conclusions
We also foresee improvements in contamination rejection and masking prior to DR1, although both of these concerns are already manageable with the Q1 data set. Because we expect that the external band photometry will improve significantly following the observation of more fields to better calibrate OU-PHZ templates, in addition to bias corrections in the photometric pipeline, we can expect that the results will improve considerably for DR1, whilst already demonstrating a successful application of multiple aspects of the Euclid cluster workflow for the detection and characterisation of optically selected galaxy clusters.
@@ Acknowledgements.
- Bulbul et al. (2024) Bulbul, E., Liu, A., Kluge, M., et al. 2024, A&A, 685, A106
@@
- Costanzi et al. (2021) Costanzi, M., Saro, A., Bocquet, S., et al. 2021, Phys. Rev. D, 103, 043522
- Cui et al. (2024) Cui, W., Jennings, F., Dave, R., Babul, A., & Gozaliasl, G. 2024, MNRAS, 534, 1247
@@
- Euclid Collaboration et al. (2025a) Euclid Collaboration, Castander, F. J., Fosalba, P., et al. 2025a, A&A, 697, A5
- Euclid Collaboration et al. (2025b) Euclid Collaboration, Ingoglia, L., Sereno, M., et al. 2025b, A&A, 695, A280
@@
- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted, Euclid Q1 SI, arXiv:2503.15302
- Euclid Collaboration: Bergamini et al. (2025) Euclid Collaboration: Bergamini, P., Meneghetti, M., Acebron, A., et al. 2025, A&A, submitted, Euclid Q1 SI, arXiv:2503.15330
- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
- Euclid Collaboration: Enia et al. (2025) Euclid Collaboration: Enia, A., Pozzetti, L., Bolzonella, M., et al. 2025, A&A, submitted, Euclid Q1 SI, arXiv:2503.15314
@@
- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H. J., Benson, K., Dolding, C., et al. 2025, A&A, submitted, Euclid Q1 SI, arXiv:2503.15303
- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
- Euclid Collaboration: Ragagnin et al. (2025) Euclid Collaboration: Ragagnin, A., Saro, A., Andreon, S., et al. 2025, A&A, 695, A282
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted, Euclid Q1 SI, arXiv:2503.15305
- Euclid Collaboration: Tucci et al. (2025) Euclid Collaboration: Tucci, M., Paltani, S., Hartley, W. G., et al. 2025, A&A, submitted, Euclid Q1 SI, arXiv:2503.15306
@@
- Lesci et al. (2022) Lesci, G. F., Marulli, F., Moscardini, L., et al. 2022, A&A, 659, A88
@@
- Romanello et al. (2025) Romanello, M., Marulli, F., Moscardini, L., et al. 2025, A&A, 693, A195
@@
- Sadibekova et al. (2024) Sadibekova, T., Arnaud, M., Pratt, G. W., Tarrío, P., & Melin, J. B. 2024, A&A, 688, A187
@@
- Tarr´ıo et al. (2019) Tarrío, P., Melin, J. B., & Arnaud, M. 2019, A&A, 626, A7
@@ Appendix B Detailed breakdown of Q1 cluster matches to external data sets
| Cluster sample | Matches with Euclid Q1 | Unique matches with Euclid Q1 |
@@
| Cluster sample | Matches with Euclid Q1 | Unique matches with Euclid Q1 |
@@
| Cluster sample | Matches with Euclid Q1 | Unique matches with Euclid Q1 |
@@ Appendix C Spectroscopic validation with OU-SPE
The detection algorithms estimate cluster redshifts based on the mean photometric redshift of member galaxies. To confirm and further refine these estimates, we examined the Euclid spectroscopy of galaxies in the cluster fields, as provided in the Q1 data release. We limited our inspections of the Euclid spectroscopy to those candidate clusters with , the lower limit at which the H lines can be detected given the spectral sensitivity range of the Euclid red grism (Euclid Collaboration: Mellier et al. 2025). This leaves us with 100 cluster candidates (17 in the EDF-F, 43 in the EDF-S, and 40 in the EDF-N).
The Euclid OU-SPE package provides five estimates for each galaxy; in the following, we only use the highest ranked estimate – spec.spec_rank=0 in the Science Archive System (SAS)– when selecting objects for verification of their spectroscopic redshifts. We restrict our analysis to galaxies with OU-SPE redshift estimates with signal-to-noise ratio (spe_line_snr_gf in the SAS) , probability (spe_z_prob in the SAS), , with line flux to avoid spurious features, and based on more than 300 pixels (spe_npix>300). Of the galaxies passing these spectroscopic quality criteria, we only considered those within 4 arcmin from the candidate cluster centre (corresponding to in the range ), and with spectroscopic redshifts within a range of , corresponding to about three times the estimated pre-launch cluster uncertainty (Euclid Collaboration: Adam et al. 2019).
In total, we selected 527 galaxies (87 in EDF-F, 197 in EDF-S, and 243 in EDF-N), and after visual inspection of their spectra we retained 170 redshift estimates as reliable (35 in EDF-F, 65 in EDF-S, and 70 in EDF-N), that is, one third of the total. This means we deal with, on average, less than two reliable redshifts per cluster field.